The contract size for platinum futures is 50 troy ounces. Suppose you need 400 troy ounces of platinum and the current futures price is $1,255 per ounce. How many contracts do you need to purchase? How much will you pay for your platinum? What is your dollar profit if platinum sells for $1,295 a troy ounce when the futures contract expires? What if the price is $1,200 at expiration? (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to the nearest whole number.)
Answer #1

Please provide rating....

1 How many contracts do you need to purchase?
Number of contract to purchase = 400/50 8 Contract
2 How much will you pay for your platinum
Payment required = 8*1255*50 = $      502,000
3 What is your dollar profit if platinum sells for $1,295 a troy ounce when the futures contract expires
Profit = =(1295-1255)*8*50 $         16,000
4 What if the price is $1,200 at expiration
Loss = =(1200-1255)*8*50 -22000